A family has 3 boys and 5 girls. In how many ways can their parents choose a basketball team of 5 people? Include evidence of your work by typing our your solution.


1
Expert's answer
2021-06-02T17:21:44-0400

The total number of children is 3 + 5 = 8

Formula for combinations is "\frac{n!}{((n-x)!\times{x!})}" , where n is a total number of permutations of a set (8), and x is a number of children in a team (5)

So the solution is "\frac{8!}{(8-5)!\times{5!}}=56"

Answer: 56 ways.